1000 nodes are connected to a 1,5 km length of coaxial cable. Using some protocol, each node can transmit 50 frames/second, where the average frame length is 100000 bits. The transmission rate at each node is 100 Mbps (where 100Gbps). What is the efficiency of this protocol?(10 pts)
Answer #1

Frame length = 100000

no.of node = 1000

Transmission rate = 100Gbps= 100×109bps

So bit transmitted by 1000 node = 50×100000×1000

= 50×108 bps

efficiency protocall = 50×108/100×109

= 0.05

% efficiency = 5%

​​​​​​